OpenStudy (anonymous):

Plutonium–238 has a yearly decay constant of 7.9 × 10-3. If an original sample has a mass of 15 grams, how long will it take to decay to 12 grams?

563blackghost (563blackghost):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\) Welcome to Openstudy! And I hope you will enjoy this helpful site! ^^ Please check out the code of conduct to learn the rules of this site xD http://openstudy.com/code-of-conduct

563blackghost (563blackghost):

Heyo! Well your question is posted in the wrong section this is biology so it would be posted in the biology or chemistry section but...here is a link to help you ^^ http://openstudy.com/updates/518d0981e4b0cf6dd8f3e29f

563blackghost (563blackghost):

whoops meant this is mathematics section ^^

OpenStudy (anonymous):

\[15e^{-.0079t}=12\] solve for \(t\)

OpenStudy (opcode):

Interestingly I forgot about how to do this stuff. (Not the original poster, just trying to refresh my memory.) An attempt: \[15e^{-.0079t}=12\] \[\dfrac{15e^{-.0079t}}{15} = \dfrac{12}{15}\] \[e^{-.0079t} = \dfrac{12}{15}\] Which do I just take \(ln()\) of both sides? If I do, why does that work? (e.g. is there a name to the property?)

OpenStudy (whpalmer4):

@opcode yes, you are on the right track. Take the natural logarithm of both sides. \[\ln(e^x) = x\]

OpenStudy (whpalmer4):

remember, the logarithm to the base \(b\) of \(u\) is some number \(a\) such that \(b^a = u\) so the log to the base of the base of an exponential just gives you the exponent. Might be easier to see with common logs. \[\log_{10} 100 = 2\]because \(10^2 = 100\) but \(100=10^2\) so \(\log_{10} (10^2) = 2\)

OpenStudy (opcode):

\[ln(e^{-.0079t}) = ln(\dfrac{12}{15})\] \[\Rightarrow .0079t = ln(\dfrac{12}{15})\] So I applied that, not sure what a natural logarithm does to a fraction, that has no exponent which in turn has no base \(e\). (Extremely rusty, I know I could just plug it into a calculator, but I am more curious on how it works.)

OpenStudy (whpalmer4):

well, a natural logarithm does the same thing with a fraction it does with any other allowable number. but there is a property of logarithms that can be used if helpful, which is that: \[\log(\frac{a}{b}) = \log(a) - \log(b)\] and for completeness: \[\log(a*b) = \log(a)+\log(b)\] (the latter being the punch line to a joke about snakes)

OpenStudy (whpalmer4):

because of that property of logs of fractions, you also have this one: \[\log(\frac{a}{b}) = \log(a) - \log(b) = -(\log(b) -\log(a)) = -\log(\frac{b}{a})\]